# Reconciling Inflation with Hubble Anisotropies

### v1

@@ 6. Example: Dirac Field as Source of Torsion
As we foresaw, this rotation is always non-trivial if the extrinsic torsion does not vanish. This “Hubble twist” is characteristic of the presence of extrinsic torsion.
@@
This gives us a dimensionless measure of the anisotropy; it is also a measure of the relative importance of extrinsic torsion to extrinsic curvature.
@@
We can see four facts very plainly, in the case of axial torsional theories of cosmic anisotropy.
[a] The predicted anisotropy singles out a distinguished direction; the Hubble parameter $H$ is the rate of separation in that direction. The rates of separation in other directions are still controlled by $H,$ so we can say that the Hubble parameter is still a measure of the overall rate of separation of the galaxies. It coincides with the “Hubble parameter” in discussions where the anisotropy is neglected, as for example in the study of the “Hubble tension”.
@@
This, then, is the predicted form of the geometric anisotropy at the end of Inflation. This geometric anisotropy would have left traces on the matter distribution at reheating. We now ask exactly how that works; the hope being that, by working backwards from current observations, one will obtain a theory that can be falsified by deductions potentially in conflict with [a], [b], [c], or [d].
@@ 7. How Torsional Inflation Sets up Initial Conditions for Reheating
here, as before, $\rho$ is the energy density, but now $S$ is the scalar curvature of the full linear connection, including spatial torsion and its derivatives, on the spatial section. Note that, in the presence of (extrinsic) torsion, $h$ is not a symmetric bilinear form, so the final term is not the square of $\hat{h}$; this proves to be important (see [37]).
@@
There are two points to notice here. The Hubble parameter itself is always isotropic, even in the presence of torsion. But we saw earlier that $\kappa$ is not necessarily an isotropic function: we decided that it is natural to assume that it is cylindrically symmetric around a distinguished direction. Equation (40) therefore tells us that the energy density at the time when reheating is about to start is also not isotropic. The anisotropy of the extrinsic geometry leaves its mark on the “initial” energy density, even though the underlying space is isotropic; and this mark may persist to recent times, as the system evolves.
@@ 8. Anisotropy and the Redshift Field
While the redshift field is no longer the sole way of observing the Hubble field, it is still the principal way. We now wish to explore how anisotropies due to extrinsic torsion affect it. (Some relevant observational references are [64, 65]; see also the discussion of these results in [22].)
The situation here is quite different to the one we contemplated in the preceding Section. There, we were concerned with anisotropies in the “initial” data, with the idea that these might ultimately be observable indirectly, in the guise of their effects on the subsequent evolution. Here, however, we are interested in direct observations of photons propagating to us from very early times, when the torsion might have been significantly larger than it is now.
@@
Let $u$ be the tangent vector to a null curve (usually called the four-dimensional wave vector), with a canonical parameter $\lambda,$ in a spacetime with torsion (not necessarily axial). In geometric optics, $u$ can be expressed as the four-dimensional gradient of a scalar eikonal function $\varphi$; that is, the one-form dual to $u$ (using the spacetime metric $g^{*}$), let us call it $U,$ can be written as $U=d\varphi.$ This is true not just for the standard Maxwell equations, but also for a very large family of generalisations of those equations [66].
@@
with respect to a coordinate basis. This is the equation for the worldlines of photons. (It is also actually the equation for the worldline of a free massive particle.)
We see that these worldlines are not, in general, geodesics. This is in sharp contrast to the case of General Relativity, where photons, as free particles, are assumed to have worldlines which are “straight lines” in spacetime.
@@
Nevertheless it seems at first rather strange that torsional theories can apparently violate something as basic as Newton’s First Law —  which is essentially what we are saying here. But this is where the study of spacetime anisotropy comes to the fore. The intuition that free particles should move along straight lines is built on the assumption that all directions in space are “equally good”: a free particle cannot “prefer” one direction over another, so it has to continue along a straight path. That would have seemed natural to Newton, but, even in ordinary General Relativity, it is very questionable, since we know that three-dimensional space can be anisotropic in general —  and this anisotropy can play a fundamental role in the physics, as it does for example in the Schwarzschild geometry.
@@
Thus, this version of the First Law rests on rather doubtful foundations even in ordinary General Relativity. In torsional theories, where, as we have seen, some kind of anisotropy is inescapable, we have even less reason to think that worldlines of free particles should be straight, so there is really no occasion for surprise when we find that they (often) are not [^18] Newton’s First Law can (and should) be salvaged by reformulating it: by declaring that a particle on which no external influence acts should have an action defined in terms of quantities pertaining only to the particle itself (its mass if any, the length or canonical parameter along its own worldline). The principle of stationary action then leads to the correct equation for the worldline, that is, to equation (44), provided that we accept that gravitation is not an “external influence”. In this way we can preserve the greatest conceptual achievement of General Relativity, the abolition of gravitational “force”..
We now turn to the question of the anisotropy of the redshift field in torsional theories (not just axial torsion theories).
Let us consider a null curve, with tangent field $u$, which is the worldline of a photon of frequency $\omega,$ as seen by the distinguished observers whose worldines are integral curves of $\xi.$ Then [67]
@@
We wish to understand how $\omega$ varies along the null curve. To this end, we compute $d\omega\over d\lambda$}, where $\lambda$ is the canonical affine parameter as above:
@@
As $u$ is a null vector, we can write it as $u=\omega\xi+\omega n,$ where $n$ is a unit spatial vector representing the direction in which the photon is moving in each spatial section through which it passes. (That is, it is a vector field along the worldline of the photon, a function of $\lambda$.) We think of its initial value as being chosen in such a manner that the photon eventually arrives at our location; so this initial vector, and therefore the subsequent values of $n$, depends on the direction in which we choose to look (though not necessarily in a simple way).
@@
where $h$ is the second fundamental form, given, in the spatially isotropic case, by equation (25) above. Since $\varkappa$ is antisymmetric, we have simply $\omega^{2}h(n,\,n)=H\omega^{2},$ since $n$ is a unit spacelike vector. Of course, this term does not depend on $n$. (Incidentally, this use of spatial isotropy marks the only point in our entire argument where we use any kind of symmetry.)
@@
where $\beta$ and $b$ are the tangential and normal components of the “bulk torsion” we defined in equations (16) and (18), above.
Combining these results, we have
${d\omega\over d\lambda}\,=\,-\,\omega^{2}\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right).$ (49)
@@
${d\omega\over d\lambda}\,=\,\omega^{2}\left(-\,H\,+\,T^{*}_{ij0}n^{i}n^{j}\,+\,T^{*}_{0i0}n^{i}\right),$ (50)
@@
The first term on the right in equation (49) is the usual one in General Relativity; it is of course isotropic. The second and third terms are not, unless the torsion takes a special form (see below). If we regard (49) as a differential equation for $\omega,$ the final value of $\omega$ when the photon is observed will be found by evaluating a solution of that equation. In general, it will of course depend on the initial value of $n$. (In fact it will also depend on the subsequent values of $n$, see below.) If so, then the redshift field will be anisotropic. It is the bulk torsion which determines whether the redshift field is isotropic.
However, something unusual occurs when the torsion is axial: $\beta(n,\,n)+b(n)=-\,T^{*}_{ij0}n^{i}n^{j}-T^{*}_{0i0}n^{i}=0.$ Thus, despite the fact that neither the extrinsic nor the bulk torsion is necessarily zero here, the redshift field for axial torsion theories is isotropic. This is consistent with the fact that the worldlines of photons are geodesics in this situation, as one sees immediately from equation (44).
@@
While this is an instructive subtlety, we ask the reader not to attach too much importance to it. We have been assuming that the torsion during the inflationary era, up to reheating, is axial; but, during reheating, the inflaton is converted to other forms of matter (but still leaves its imprint on those other forms). To be concrete, let us assume that the immediate post-inflationary matter can be represented as a fluid. As this fluid will inherit the intrinsic spin of the inflaton, it can be represented (very) approximately by a Weyssenhoff fluid (see for example [68]). The torsion corresponding (via the Einstein-Cartan field equations) to such a fluid is not axial.
In more detail: the spin tensor of a Weyssenhoff fluid takes the form $v\,\otimes\,\sigma,$ where $v$ is the unit tangent vector field to the worldlines of the fluid, and $\sigma$ is a two-form representing the spin, which is taken to reside in a spatial section defined by $v$ (meaning that the spacetime one-form $\sigma(v,\,\_)$ vanishes identically). We will assume, as above, that the initial state of the fluid is dictated by the state of the inflaton at the start of reheating: so it inherits the geometric structure we discussed above. That is, it will be cylindrically symmetric (as seen from a distinguished spatial section) around a distinguished axis. It is reasonable to surmise that this distinguished axis leaves its mark on the Weyssenhoff fluid in the form of the eigenvector of $\sigma$ that has zero eigenvalue. (This eigenvector must exist, for the same reason as in our discussion in Section 6.)
In the Einstein-Cartan context, the torsion is then
@@
In general, then, the fluid worldlines are not perpendicular to the distinguished spatial sections $\Sigma$ in which the Universe is homogeneous. One says that such a cosmology is “tilted” [69]. Tilted cosmologies have recently seen a revival: see [70] for a particularly interesting example, and [71] for stimulating observations on this question.
@@
To summarise, then: the redshift field is, in a torsional theory, almost certainly anisotropic after reheating. But the photons we observe do not emanate from matter during that period: they come from the later time of decoupling (or “recombination”). In the interim, the photons cannot stream freely, so it may well be that all signs of the immediate post-Inflationary anisotropy have been lost by the time decoupling occurs. So we have to ask: can any torsion which survives to the time immediately after decoupling give rise to still further anisotropy, which might be directly observable?
@@
By the time of decoupling/recombination, when the CMB photons began their journey towards us, we can assume that the spacetime metric is approximately of the general FRW form, with Euclidean spatial sections and a scale factor $a(t)$, where $t$ is the usual proper cosmic time. Let us suppose that the torsion, in particular the bulk torsion, is not yet negligible at this time and for an interval afterwards. (We will not assume that it arises from a Weyssenhoff fluid at this relatively late time, when the fluid approximation has broken down; instead we return to considering torsion of an arbitrary form.)
@@
The cosmological redshift is often explained in the following elementary way. When the photon arrives at our location, its wavelength has been “stretched” by the cosmic expansion, by a factor of $a_{\mathrm{final}}\over a_{\mathrm{initial}}.$ Therefore its frequency has been reduced by that same factor, and so we define a redshift factor $z$ in the usual way, $z\,=\,{\omega_{\mathrm{initial}}\over\omega_{\mathrm{final}}}-1\;=\;{a_{\mathrm{final}}\over a_{\mathrm{initial}}}-1.$ If this is so, if the final frequency depends only on the initial and final values of the scale factor, then it seems that no further anisotropy can arise.
But is this actually the case? Is it clear that whatever conditions the photon experiences, in the vast “bulk” of spacetime from decoupling to our location, have no effect on the redshift? We need to investigate [^20] The following argument is essentially the solution of the suitably generalised version of Problem 4 in Chapter 5 of [67]..
@@
Since $H=(da/dt)/a$ in this case, and $\omega=dt/d\lambda,$ we have $a\omega H=da/d\lambda,$ so that there is a cancellation and we obtain
@@
If we set both parts of the bulk torsion equal to zero, we obtain a remarkable result: $a\omega$ has the same value when the photon is received as it had initially. This validates the above elementary explanation of the redshift: at the point of observation, the wavelength of the photon has been “stretched” by a factor $a_{\mathrm{final}}/a_{\mathrm{initial}}$, so the frequency (and the energy) has been reduced accordingly, because $a\omega$ is constant along the worldline.
But what is the physics underlying the constancy of $a\omega?$ Let us try to explain this.
In a time-dependent spacetime geometry, there is no timelike Killing vector field, and consequently no law of conservation of energy. The photon simply loses energy at some points along its path, and gains it at others, depending on the way in which the geometry varies with time; and there is no reason to expect that the gains at one time will compensate for the losses at another.
If however we imagine for example a zero-torsion cosmology with a periodically fluctuating scale factor, then the fact that $a\omega$ is constant means that the photon always does regain, over a complete cycle, whatever energy it has lost: there is no “memory” of any losses or gains. This is the real reason for the fact that the final energy loss depends only on the ratio $a_{\mathrm{final}}/a_{\mathrm{initial}}$ in the usual argument. The absence of what might be called “cosmic hysteresis” in zero-torsion cosmology is the real physical meaning of the constancy of $a\omega.$
When the bulk torsion makes its presence felt, however, the situation is very different. When equation (56) is integrated, we obtain
@@
so we find that the ratio of the final frequency to the initial frequency depends on the values taken by the bulk torsion, and on the values of the spatial direction field $n$ (and also on the frequency itself) all along the path —  not just on the initial and final values of the scale factor. When the bulk torsion is not zero, the final redshift can no longer be understood simply in terms of the stretching of wavelengths: the loss of energy depends on the geometry of the far-off and long-ago “bulk” of spacetime as the photon passes through it. In short, bulk torsion gives rise to “cosmic hysteresis”.
The dependence on $n$, and on the bulk torsion fields if (as is generically the case) they themselves are not isotropic, means that the redshift field in a torsional theory will be anisotropic in general.
Even if the bulk torsion is negligible currently and was so in the recent past, the hysteresis could still have a potentially detectable anisotropic effect on the final redshift if the bulk torsion was not negligible even for a brief period in the distant past. Specifically: if there was an appreciable bulk torsion at and shortly after decoupling, then equation (57) means that it might have had an observable effect on the redshift now, even if the bulk torsion has diluted, to the point of being unobservable, since that time.
The exact form taken by this anisotropy is difficult to predict, precisely because of the dependence on the values of $\beta,$ $b$, $\omega$, and $n$ along the worldline of the photon; the integral on the right side of equation (56) will be very difficult to evaluate in practice. No doubt a numerical investigation would be necessary to determine it. But the essential point is clear: in a torsional theory, the redshift field may well be slightly anisotropic.
We noted, in Section 1, that we proposed to leave to one side other, non-geometric, potential sources of cosmic anisotropy, including the possibility of bulk flows [34, 35]. We have seen, in our discussion of the Weyssenhoff fluid, that “tilt” arises very naturally in torsional theories, so the two kinds of anisotropy might well occur simultaneously and be intertwined. It might be challenging to separate the two effects.

### v2

@@ 6. Example: Dirac Field as Source of Torsion
As we foresaw, this rotation is always non-trivial if the extrinsic torsion does not vanish. This “Hubble twist” is particularly simple in the case of axial torsion.
@@
This gives us a dimensionless measure of the anisotropy; it is also a measure of the relative importance of extrinsic torsion compared to extrinsic curvature.
@@
We can see four facts very plainly, in the case of axial torsional theories of the geometry of spacetime during Inflation.
[a] The predicted anisotropy singles out a distinguished direction; the Hubble parameter $H$ is the rate of separation in that direction. The rates of separation in other directions are still controlled by $H,$ so we can say that the Hubble parameter is still a measure of the overall rate of separation of the galaxies. It plays the role of the “Hubble parameter” in discussions where the anisotropy is neglected, as for example in the study of the “Hubble tension”.
@@
This, then, is the predicted form of the geometric anisotropy at the end of Inflation. This geometric anisotropy might have left traces on the matter distribution at reheating. We now ask exactly how that works; the hope being that, by working backwards from current observations, one will obtain a theory that can be falsified by deductions potentially in conflict with [a], [b], [c], or [d]. (It is important to notice, however, that some observations of cosmic anisotropy are concerned with spacetime geometry long after the end of Inflation, so some of these remarks may not apply to them. This is particularly the case for the redshift field, as we will discuss below in Section 8.)
@@ 7. How Torsional Inflation Sets up Initial Conditions for Reheating
here, as before, $\rho$ is the energy density, but now $S$ is the scalar curvature of the full linear connection, including spatial torsion and its derivatives, on the spatial section. Note that, in the presence of (extrinsic) torsion, $h$ is not a symmetric bilinear form, so $\hat{h}^{T}\hat{h}$ is not the square of $\hat{h}$; this proves to be important (see [52]).
@@
There are two points to notice here. The Hubble parameter itself is always isotropic, even in the presence of torsion. But we saw earlier that $\kappa$ is not necessarily an isotropic function: we decided that it is natural to assume that it is cylindrically symmetric around a distinguished direction. Equation (40) therefore tells us that the energy density at the time when reheating is about to start is also not isotropic. The anisotropy of the extrinsic geometry leaves its mark on the “initial” energy density, even though the underlying space is isotropic; and this mark may persist to later times, as the system evolves.
@@ 8. Anisotropy and the Redshift Field
While the redshift field is no longer the sole way of observing the Hubble field, it is still the principal way. We now wish to explore how anisotropies due to torsion affect it.
The situation here is quite different to the one we contemplated in the preceding Section. There, we were concerned with anisotropies in the “initial” data, with the idea that these might ultimately be observable indirectly, in the guise of their effects on the subsequent evolution. Here, however, we are interested in direct observations of photons propagating to us from various early times, when the torsion might have been significantly larger than it is now.
@@
Let $u$ be the tangent vector, usually called the four-dimensional wave vector, to a null curve with a parameter $\lambda$ (which we can modify if necessary) in a spacetime with torsion (of any kind). In geometric optics, $u$ can be expressed as the four-dimensional gradient of a scalar eikonal function $\varphi$; that is, the one-form dual to $u$ (using the spacetime metric $g^{*}$), let us call it $U,$ can be written as $U=d\varphi.$ This is true not just for the standard Maxwell equations, but also for a very large family of generalisations of those equations [80].
@@
with respect to a coordinate basis, where of course the coefficients $T^{*c}_{ab}$ are the components of the spacetime torsion tensor. This is the equation for the worldlines of photons. (It is also actually the equation for the worldline of a free massive particle.)
We see that these worldlines are not, in general, geodesics (though they can be). This is in sharp contrast to the case of General Relativity, where photons, as free particles, are assumed to have worldlines which are always “straight lines” in spacetime.
@@
Nevertheless it seems at first rather strange that torsional theories can apparently violate something as basic as Newton’s First Law (“A free particle moves along a straight line at constant speed.”) —  which is essentially what we are saying here. But this is where the study of anisotropy comes to the fore. The intuition that free particles should have straight worldlines is built on the assumption that space and time are (in a sense that can readily be made precise) uniform: a free particle cannot in that case “justify” deviating from straight line motion, or moving at a non-constant speed. This assumption of uniformity seemed natural to Newton, and it also holds true in Special Relativity; but, even in ordinary General Relativity, it does not. For we know that, in that theory, three-dimensional space is not “uniform” in general —  and this can play a fundamental role in the physics, as it does for example in the Schwarzschild geometry. Similarly, time is not “uniform” in spacetimes, such as cosmological spacetimes, that lack a timelike Killing vector field.
@@
Thus, this version of the First Law rests on very doubtful foundations even in ordinary General Relativity. In torsional theories, where, as we have seen, some kind of spatial anisotropy is almost inescapable, we have even less reason to think that worldlines of free particles should be straight, so there is really no occasion for surprise when we find that they (often) are not [^18] Newton’s First Law can (and should) be salvaged by reformulating it: by declaring that a particle on which no external influence acts should have an action defined in terms of quantities pertaining only to the particle itself (its mass if any, the length or canonical parameter along its own worldline). The principle of stationary action then leads to the correct equation for the worldline, that is, to equation (44), provided that we accept that gravitation is not an “external influence”. In this way we can preserve the greatest conceptual achievement of General Relativity, the abolition of gravitational “force”..
In the special case of “isotropic torsion”, the right side of equation (44) is $(\phi\xi^{a}u_{a})u_{b}$ (see [67]), which is of the form (function of the parameter $\lambda$) $\times\,u_{b}$. In this case, it is well known that a change of parameter reduces the right side of (44) to zero, so the worldline of a photon in the isotropic case is in fact a geodesic. This underlines the claim that the worldlines of photons are usually not geodesics in torsional theories, because such theories usually involve some kind of anisotropy [^19] The precise statement is that when a torsional spacetime is isotropic, then the worldlines of particles are geodesics. The converse is not true, because sometimes anisotropy is present but the particles may not be able to detect it. This happens, in fact, when the torsion is axial: see below..
We now turn to the question of the anisotropy of the redshift field in torsional theories.
Let us consider a null curve, with tangent field $u$, which is the worldline of a photon of frequency $\omega,$ as seen by the distinguished observers whose worldines are integral curves of $\xi.$ As $u$ is a null vector, and because the timelike component of $u$ is the frequency [81], we can write it as $u=\omega\xi+\omega n,$ where $n$ is a unit spatial vector representing the direction in which the photon is moving in each spatial section through which it passes. (That is, it is a vector field along the worldline of the photon, a function of $\lambda$, just as $\omega$ is a scalar function of $\lambda$.) We think of its initial value as being chosen in such a manner that the photon eventually arrives at our location; so this initial vector can be regarded as dependent on the direction in which we choose to look (though not necessarily in a simple way).
Equation (43) (or (44)) is a differential equation for $u$, but we can think of it as a pair of simultaneous ordinary differential equations for $\omega$ and $n$, as functions of $\lambda$. Normally the variables in such a pair of non-linear, intertwined equations cannot be separated; if that is so, the final value of $\omega$ depends not only on the initial value of $\omega$ but also on the initial value of $n$. In that case, the final, observed value of $\omega$ depends on the direction in which we choose to look: in other words, the redshift field is anisotropic.
Let us see how this works in detail. First, we must find the equation for $\omega$ explicitly.
We have
@@
We wish to compute $d\omega\over d\lambda$}, where $\lambda$ is the canonical affine parameter as above:
@@
where $h$ is the second fundamental form. (The appearance of $h$ here is how the Hubble field will enter the discussion.)
@@
where $\beta$ and $b$ are the tangential and normal components of the “bulk” torsion we defined in equations (16) and (18), above.
Combining these results, we have (simplifying by using $\omega^{-1}$ instead of $\omega$)
${d\omega^{-1}\over d\lambda}\,=\,h(n,\,n)\,+\,\beta(n,\,n)\,+\,b(n).$ (49)
@@
${d\omega^{-1}\over d\lambda}\,=\,h_{ij}n^{i}n^{j}\,+\,T^{*}_{i0j}n^{i}n^{j}\,+\,T^{*}_{00i}n^{i},$ (50)
@@
Clearly, in general, $n$ is very much present on the right side of equation (49), and $\omega$ cannot be separated from it; and so, integrating this equation, we will find that the final value of $\omega$, that is, the redshift field, is generically anisotropic, especially if “bulk” torsion is present. In particular, $b$, being essentially a spatial vector field, cannot be isotropic if it is not zero.
We have said that anisotropy corresponds to a dependence of the final value of $\omega$ on the “initial” value of $n$. But $n$ is not a constant along the worldline in general, so this “initial” value will depend on our choice as to when the “initial” time should be. This means that the extent of the anisotropy in the redshift field can and usually will be time-dependent in the presence of torsion. We will return to this important observation at the end of this Section.
If we restrict to a strictly metric theory, such as General Relativity, then the torsion terms on the right side of equation (49) are zero. If we assume that Inflation performs its customary service of isotropising the spatial sections, then, as usual, the second fundamental form inherits isotropy, and this is the essential point: we have $h=Hg$, and so, since $n$ is a unit vector, this gives us h(n,n)=H,h(n,\,n)=H, where now $H$ is to be regarded as a function of $\lambda$ (only). Then (49) is greatly simplified, and when integrated it gives
$\omega^{-1}\,=\,\int\,H\,d\lambda.$ (51)
All dependence on $n$ has dropped out; the change in the value of $\omega$ is due exclusively to the cumulative effect of the “stretching” described by the Hubble parameter, which is inherently isotropic. It follows that, in conventional Inflationary theories embedded in General Relativity or another strictly metric theory, the redshift field inherits isotropy from the spatial geometry, in the familiar manner. (A similar argument works in the case of “isotropic torsion” [67]: again, the redshift field is always isotropic in that theory if the intrinsic spatial geometry is so, since $b(n)=0$ and $\beta(n,\,n)=4\phi g,$ so the integrand in equation (49) is independent of $n$: it is equal to $H\,+\,2\phi,$ and equation (51) is valid if modified in the same way.)
Thus, in strictly metric theories of gravitation, redshift anisotropy can only be present if it occurs in the earliest relevant times, that is, in the CMB; but, again, that would be hard to reconcile with Inflation. Furthermore, such anisotropies would tend to dissipate with the passing of time [12] (but see the potentially very important novel viewpoint on this put forward in [82]).
This is presumably why, when, in [29], substantial anisotropies were reported, it was claimed that these observations had “already been refuted” by the fact that anisotropies at that level (or higher) were not seen in the CMB. We will argue that this refutation was indeed correct, but only if torsion is neglected.
Returning to the torsional case, we have, when the intrinsic geometry of the spatial sections is isotropic,
$\omega^{-1}\,=\,\int\,\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right)\,d\lambda.$ (52)
We stress again that the right side [^20] We notice in passing that the extrinsic torsion plays no role here : it drops out because the extrinsic torsion form is antisymmetric (see equation (22)). So it is still the case that h(n,n)=H.h(n,\,n)=H. Instead the kind of torsion responsible for redshift anisotropy is the “bulk” torsion. tells us both that the redshift field is generically anisotropic in the presence of torsion, and that the “anisotropy field” varies with time.
However, something unusual occurs when the torsion is axial : $\beta(n,\,n)+b(n)=T^{*}_{i0j}n^{i}n^{j}\,+\,T^{*}_{00i}n^{i}=0:$ torsion is present, but photons are insensitive to it. Thus, despite the fact that neither the extrinsic nor the bulk torsion is necessarily zero here, the redshift field for axial torsion theories is isotropic in a given epoch if the intrinsic spatial geometry is isotropic then.
@@
While this is an instructive subtlety, we ask the reader not to attach too much importance to it. We have been assuming that the torsion during the inflationary era, up to reheating, is axial; but, during reheating, the inflaton is converted to other forms of matter (but still leaves its imprint on those other forms). To be concrete, let us assume that the immediate post-inflationary matter can be represented as a fluid. As this fluid will inherit the intrinsic spin of the inflaton, it can be represented approximately by a Weyssenhoff fluid (see for example [83]). The torsion corresponding (via the Einstein-Cartan field equations) to such a fluid is not axial.
In more detail : the spin tensor of a Weyssenhoff fluid takes the form $v\,\otimes\,\sigma,$ where $v$ is the unit tangent vector field to the worldlines of the fluid, and $\sigma$ is a two-form representing the spin, which is taken to reside in a spatial section defined by $v$ (meaning that the spacetime one-form $\sigma(v,\,\_)$ vanishes identically). In the Einstein-Cartan context, the torsion is then
@@
We will assume, as above, that the initial state of the fluid is dictated by the state of the inflaton at the start of reheating : so it inherits the geometric structure we discussed above. In fact, $\sigma$ is just a non-zero multiple of the extrinsic torsion form at this time, and it follows that its non-zero eigenvalue will be cylindrically symmetric (as seen from a distinguished spatial section) around a distinguished axis, defined by the eigenvector of $\sigma$ that has zero eigenvalue.
@@
In general, then, the fluid worldlines are not perpendicular to the distinguished spatial sections $\Sigma$ in which the Universe is homogeneous. One says that such a cosmology is “tilted” [84]. Tilted cosmologies have recently seen a revival : see [82] for a particularly interesting example, and [85] for stimulating observations on related questions.
@@
In the context of standard General Relativity, tilted cosmologies can account for anisotropies in the observations [82], but this is achieved at the expense of having anisotropic spatial sections; as we know, this cannot be avoided in that case. The advantage of the torsional approach is that one can obtain the same objective without spatial anisotropy, in accord with Inflation.
To summarise, then : the redshift field is, in a torsional theory, almost certainly anisotropic after reheating.
@@
But the photons we observe do not emanate from matter during that period : they begin to come from the later time of decoupling (or “recombination”). In the interim, the photons cannot stream freely, so it may well be that all signs of the immediate post-Inflationary anisotropy in the redshift field have been lost by the time decoupling occurs. So we have to ask : can any torsion which survives to the time immediately after decoupling give rise to still further redshift anisotropy, which might be directly observable?
@@
By the time of decoupling/recombination, when the CMB photons began their journey towards us, we can assume that the spacetime metric is approximately of the FRW form, with Euclidean spatial sections and a scale factor $a(t)$, where $t$ is the cosmic time defined by using Gaussian normal coordinates constructed using $\nabla^{0},$ the zero-torsion connection defined by $g^{*}$. (Then $\xi=\partial_{t}$ when both are evaluated on $\Sigma;$ this relation need not hold off $\Sigma,$ but that will not affect our discussion here.) Let us suppose that the torsion, in particular the bulk torsion, is not yet negligible at this time and for an interval afterwards. (We will not assume that it arises from a Weyssenhoff fluid at this time; instead we return to considering torsion of an arbitrary form.)
We now ask: how does torsion affect the redshift field we actually observe?
@@
In conventional General Relativistic cosmology, one often sees an argument of the following form. When the photon arrives at our location, its wavelength has been “stretched” by the cosmic expansion, by a factor of $a_{\mathrm{final}}\over a_{\mathrm{initial}}.$ Therefore its frequency has been reduced by that same factor, and so if we define a redshift factor $z$ in the usual way, $z\,=\,{\omega_{\mathrm{initial}}\over\omega_{\mathrm{final}}}-1$, then $z$ is given simply by $z\;=\;{a_{\mathrm{final}}\over a_{\mathrm{initial}}}-1.$
This argument is however based on the assumption that “stretching” is the only effect to which the wavelength is subjected. In the presence of “bulk” torsion, that assumption is not correct. Separately, the argument is also based on the assumption that the history of the photons does not matter.
We can explain this second point in a more physical way as follows. In a time-dependent spacetime geometry, there is no timelike Killing vector field, and consequently no law of conservation of energy. The photon simply loses energy at some points along its path, and gains it at others, depending on the way in which the spatial geometry varies with time; and there is no reason to expect that the gains at one time will compensate for the losses at another. Yet the elementary argument we have been discussing insists that, if for example $a_{\mathrm{final}}$ happens to be the same as $a_{\mathrm{initial}}$, then this compensation must happen and be complete and exact. One could say that the argument is equivalent to the claim that there is no “memory” of the losses and gains of energy, no “cosmic hysteresis”.
As we will see shortly, this is actually correct (in General Relativity), but this way of putting the point makes it clear that the standard “elementary” argument is far more subtle than it appears.
Clearly we need a better understanding of the whole argument [^22] The following discussion is essentially the solution of the suitably generalised version of Problem 4 in Chapter 5 of [81]..
@@
Since $H=(da/dt)/a$ here, and $\omega=dt/d\lambda,$ we have $a\omega H=da/d\lambda,$ so that there is a cancellation and we obtain
@@
If we set both parts of the bulk torsion equal to zero, we obtain a remarkable result : $a\omega$ is constant along the worldline, so it has the same value when the photon is received as it had initially. This validates the above elementary explanation of the redshift  in General Relativity: at the point of observation, the wavelength of the photon has indeed been “stretched” by a factor $a_{\mathrm{final}}/a_{\mathrm{initial}}$, so the frequency (and the energy) has been reduced accordingly, without regard to anything that the photon encountered along its worldline; there is no “hysteresis”. But this is only true because $a\omega$ is constant along the worldline.
The tale is very different when “bulk” torsion is present, however. Then, when equation (58) is integrated, we obtain
@@
so we find that the ratio of the final frequency to the initial frequency depends on the values taken by the “bulk” torsion all along the path (because $\beta$ and $b$ are evaluated on $n$) —  not just on the initial and final values of the scale factor. When the bulk torsion is not zero, the final redshift can no longer be understood simply in terms of the stretching of wavelengths : the loss of energy depends on the geometry of the far-off and long-ago “bulk” of spacetime as the photon passes through it, and this effect could vary for photons passing through different regions of spacetime —  it could be path-dependent. In short, bulk torsion gives rise to “cosmic hysteresis”.
In torsional inflationary cosmologies, then, the observed redshift field is generically anisotropic, and the extent of the anisotropy depends on time. Of course, the manner in which the redshift anisotropy varies with time will depend on the particular theory chosen. One can speculate that, in some cases, the “anisotropy field” could fluctuate upwards, perhaps quasi-periodically, before eventually dying out. In such a theory one might not be able to argue against claims of observed anisotropy at later times by using the absence of anisotropy at that level in the CMB, and this could open the way to a full theory for observed anisotropies, if they exist.